OpenStudy (anonymous):

a 25 inch t.v. set also has a width of 20 inches what is the height of the t.v. set?

OpenStudy (anonymous):

I got 15 inches in height am I right

OpenStudy (anonymous):

Height = sqrt(25^2 - 20^2) = sqrt(625 - 400) = sqrt(225) = 15

Directrix (directrix):

You can use the Pythagorean Theorem: 25^2 = 20^2 + h^2 to solve and get 15. I used Pythagorean Triples. Have you studied those yet. Big time saver.